Calculate mass balance of 2000 kg sulfuric acid production from gaseous sulfur dioxide in the air conditions. 84% of water is used for production
1
Expert's answer
2020-06-18T14:46:02-0400

Reactions of industrial production of sulfuric acid:

S+O2=SO2

2SO2+O2=2SO3

SO3+H2O=H2SO4

It follows that 1mol of SO2 accounts for 0.5mol of O2, 1mol of H2O and 1mol of H2SO4.

nH2SO4=mH2SO4/MH2SO4=2000*103g/98g*mol-1=20408mol

nH2SO4=nH2O=nSO2=20408mol

nO2=0.5*nH2SO4=10204mol

mH2O=20408mol*18g*mol-1=367344g=367.344kg

mSO2=20408mol*64g*mol-1=1306112g=1306.112kg

mO2=10204mol*32g*mol-1=326528g=326.528kg

Since only 84% of water and therefore only 84% of oxigen and sulfur dioxide are used in the production of acid, we will get the material balance of production of 2000kg of sulfuric acid:

For 2000kg of H2SO4 we need 367.344kg/0.84=437.314kg≈437kg of H2O

1306.112kg/0.84=1554.895kg≈1555kg of SO2

326.528kg/0.84=388.724kg≈389kg of O2

Answer:

For 2000kg of H2SO4 we need 437kg of H2O

1555kg of SO2

389kg of O2